# SPECIAL RELATIVITY

### v11

@@ 12.5 Metrics in Accelerated Frames
Let us examine the implications of assuming that the inertial frame has no history of acceleration and that the space-time metric, as perceived by an observer at rest, is diagonal. We begin by considering two emitters within this inertial frame: the first is stationary, while the second moves with velocity $v$. Applying a Galilean boost yields the metric of the accelerated emitter in the inertial frame, as given in Eq. (11). In this coordinate system, the non-accelerated light source is described by the standard diagonal metric: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
Let us now return to the discussion of the accelerated frame. In this frame, we observe a reciprocal velocity: the second emitter is at rest, while the first emitter moves with velocity $-v$. Now what about reciprocity in physics? According to conventional theory, acceleration does not inherently break the symmetry of motion between an accelerated frame and an inertial one. Consequently, an observer in the accelerated frame should still obtain a diagonal metric for the second emitter, which is at rest relative to this frame. By the principle of reciprocity, the metric given in Eq. (15) should also describe the aberration of light as observed for the first emitter, which appears to be moving in the accelerated frame. In fact, the metric of the moving emitter in the accelerated frame matches that of the moving emitter in the inertial frame, as given by Eq. (11). Reciprocity implies that the only difference lies in the sign of the velocity $v$. The principle of reciprocity states that all inertial frames are equivalent. However, this relativity principle does not hold and the Lorentz covariance of all physical equations does not dictate the reciprocal symmetry of nature.
We will examine the asymmetry between inertial and accelerated reference frames. A space-time geometrical analysis of the accelerated frame reveals a fundamental non-reciprocity between it and the initial inertial frame. Specifically, within the accelerated frame, we derive the Langevin metric, Eq. (15), associated with the second emitter. This metric governs the behavior of electrodynamics in the accelerated frame. Our analysis highlights a crucial asymmetry: Maxwell’s equations, which hold in inertial frames, are no longer valid from the perspective of an observer at rest in the accelerated frame. This breakdown reflects a deeper distinction between the two types of frames. Importantly, because we adopt an absolute time coordinatization, the initial conditions—such as the orientation of radiation wavefronts—remain identical in both frames  [^45] Many physicists contend that the universality of Maxwell’s equations across inertial frames rules out the existence of any privileged frame. As Dieks observes [51], “As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.” However, this widely held view is misleading. The metric alone does not fully specify the physical situation. To solve electrodynamics problems, one must also provide appropriate initial conditions..
We now turn to an apparent paradox. The principle that all inertial frames are equivalent is often supported by the following argument: To maintain Lorentz coordinates in an accelerated frame, one must perform a diagonalization of the metric tensor. In both the original and the accelerated frames, the metric appears diagonal, and standard treatments suggest that the coordinates in these frames should be related by a Lorentz transformation. At first glance, the situation appears symmetric—both frames seem equivalent. However, a crucial asymmetry emerges when we consider the physical distinction between an inertial frame and one that has undergone acceleration. Specifically, we should expect the metric tensor to reflect this difference. After the diagonalization procedure, the metric tensor in the accelerated frame must transition discontinuously—from a form such as $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the standard inertial form $g_{00}=1,g_{01}=0$. This discontinuity implies that the Lorentz transformation does not preserve the smooth structure of the metric across frames. As a result, the symmetry between inertial frames breaks down in this context. This reveals a subtle yet significant distinction: while the mathematical form of the metric may appear similar in both frames, the physical histories of the frames—particularly the presence or absence of acceleration—lead to observable and meaningful differences. In this way, the accelerated frame retains a kind of “memory” of its non-inertial past, breaking the naive equivalence suggested by symmetry arguments.
@@ 13.1 The Concept of Time in Special Relativity
Here is a simple example that illustrates the concept clearly. Suppose we know the law governing muon decay in its rest frame. When we apply a Lorentz transformation to this decay law, we find that, in the laboratory frame, the characteristic lifetime of the particle appears to increase from $\tau_{0}$ to $\gamma\tau_{0}$. This result can be interpreted as follows: after traveling a distance of $\gamma v\tau_{0}$, the number of muons observed in the lab frame is reduced to half of the original population. In this context, the time measured corresponds to a ”length,” as defined by the earlier metric discussion. This leads to an intriguing question: could this interpretation apply to all time measurements within the framework of special relativity? We argue that the answer is yes.
@@ 13.2 Interference Phenomena
Suppose the laboratory observer measures the wavelength of a standing wave. The relation $dl_{i}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ describes the connection between the spatial line element and the metric. Using Eq. (11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which defines the spatial geometry in the context of the Langevin metric. This result reflects the Lorentz contraction of measuring rods. When analyzing the measured data, the laboratory observer finds that the wave number is given by $\sqrt{1-v^{2}/c^{2}}\omega_{0}/[c(1-v/c)]=\omega_{0}\sqrt{1+v/c}/[c\sqrt{1-v/c}]$. Notably, this is the same factor one obtains under the assumption of a diagonal metric. In other words, the laboratory observer will measure the same two-way speed of light regardless of the choice of metric representation.
@@ 14.1 Second-Order Optical Experiments and Special Relativity
An illustrative example of time dilation is observed in the resonance absorption of gamma rays in Mössbauer rotating disk experiments. When measuring the resonance absorption for the same relative velocity between the source and the observer, a blue shift is detected when the source is positioned at the center of the disk, while a red shift appears when the source is located at the rim. This result reveals an asymmetry in the transverse Doppler effect [^46] The first experimental confirmation of time dilation via resonance absorption of gamma rays was reported by Champeney, Isaak, and Khan in 1965 [52].. Notably, when both the source and observer are situated at the rim, no frequency shift is observed. In such a configuration, a pulse emitted by one accelerated observer on the rotating disk, which appears redshifted at the center due to time dilation, is received by a diametrically opposite observer on the rim without any frequency change. This occurs despite both observers moving at the same speed $v$ in opposite directions in the laboratory frame. Their mutual time dilation effects effectively cancel the frequency shift in this scenario.
@@ 14.3 Wigner Rotation and the Trouton-Noble Experiment
Consider two opposite point charges $+e$ and $-e$ in the Earth-based frame; the radius-vector pointing from $-e$ to $+e$ be denoted by $\vec{r}_{n}$. If the charges are at rest the force acting upon $+e$ can be written $\vec{F}_{n}=e\vec{E}=-e^{2}\vec{r}_{n}/r^{3}_{n}$. As the force acts in the direction $\vec{r}_{n}$ the moment of force produced by the pair of charges vanishes, $\vec{M}_{n}=\vec{r}_{n}\times\vec{F}_{n}=0$.
If the pair of charges is made to move with a velocity $\vec{v}$ in the Sun-centered frame then the positive charge will be under the action of the Coulomb attraction of $-e$ and also under the influence of the magnetic field $\vec{B}=e(\vec{v}\times\vec{r})/(cr^{3})$. Thus the total force acting upon $e$ is given by $\vec{F}+e^{2}(\vec{v}\times(\vec{v}\times\vec{r}))/(c^{2}r^{3})$. Since $\vec{v}\times(\vec{v}\times\vec{r})=\vec{v}(\vec{v}\cdot\vec{r})-v^{2}\vec{r}$ we find that the moment of force produced by the pair of charges is equal to $\vec{M}=e^{2}(\vec{v}\cdot\vec{r})(\vec{r}\times\vec{v})/(c^{2}r^{3})$. Denoting the angle between $\vec{v}$ and $\vec{r}$ by $\theta$, we find for the absolute value of moment of force $M=e^{2}v^{2}\sin{\theta}\cos{\theta}/(c^{2}r)$.
@@
The absence of torque on the capacitor in the Earth-based frame is fully consistent with special relativity and can be explained in a straightforward manner. The electrodynamic equations in the Earth-based frame can be derived by applying a Galilean transformation (with velocity $-v$) to Maxwell’s equations. Assuming the magnetic field $\vec{B}=0$ in the original frame, the transformed fields are given by $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$, where $\vec{v}$ is the velocity of the Earth relative to the Sun-based frame. Consequently, the induced magnetic field in the Earth-based frame does not exert any torque on the capacitor, which is at rest in that frame.
The presence of a magnetic field gives rise to a magnetic force, which, in turn, exerts a torque on a charge in the sun-based inertial frame. This leads to an apparent paradox: the mechanical equations governing force and torque appear to differ between inertial frames. Specifically, a torque—and thus a time-dependent change in three-dimensional angular momentum—exists in one inertial frame but not in another that is in relative motion  [^48] In [55], Jackson discusses the apparent paradox that different mechanical equations for force and torque govern the motion of a charged particle in different inertial frames. A similar paradox arises in the standard explanation of the Trouton–Noble experiment. In the usual approach [55], this paradox stems from the hidden assumption that all inertial frames share a common three-dimensional space. However, we demonstrate that a relativistically accurate description of the Trouton–Noble experiment can be achieved only by taking into account the Wigner rotation effect.. We argue that the root of this paradox lies in a flawed assumption: that an Earth-based observer and a Sun-based observer share a common three-dimensional space. According to special relativity, observers following different worldlines (i.e., with different trajectories through spacetime) possess different three-dimensional hypersurfaces of simultaneity—different 3-spaces. The standard treatment of the Trouton-Noble experiment implicitly assumes that the coordinate axes $(x_{n},y_{n},z_{n})$ of the moving Earth-based observer are parallel to the axes $(x,y,z)$ of the stationary Sun-based observer. We demonstrate that a consistent explanation, fully aligned with special relativity, emerges only when the equations of motion are expressed in covariant form. This approach properly accounts for the relativistic transformation of coordinate systems and resolves the apparent contradiction.
@@ 14.4 Time Dilation in Moving Atomic Clock
In 1971, Hafele and Keating conducted a landmark experiment in which four atomic clocks were synchronized with a laboratory reference clock and then flown around the Earth aboard commercial aircraft—two traveling eastward and two westward  [^49] This experiment provided the first experimental confirmation of time dilation as predicted by special relativity, using atomic clocks in circular motion. An important follow-up came from the Global Positioning System (GPS), which also involves relativistic effects. GPS satellites, numbering around 80, orbit the Earth at velocities of approximately 4 km/s. The system demonstrates that satellite clocks tick more slowly due to their relative motion, with a measured discrepancy of about 7.1 microseconds per day. Interestingly, while there is a noticeable time difference between the clocks on Earth and those in orbit, no asymmetries are observed among the satellites themselves. This suggests that relative velocity alone does not account for time dilation. Furthermore, the experiment revealed no effect of Earth’s orbital velocity on clock rates, a result attributed to the fact that both the satellites and Earth-based clocks share a common orbital path around the Sun, rendering orbital effects unobservable in this context [56, 57].. Upon return, the airborne clocks were compared with the laboratory reference clock. The results showed that the clocks carried on the planes had run slower than those left stationary on Earth. The observed discrepancies matched the predictions of special relativity, confirming the role of relative acceleration in time dilation. Notably, the experiment found no measurable influence of Earth’s orbital motion on clock rates, reinforcing the idea that time dilation effects are tied to local motion relative to the chosen inertial frame.
[…]
@@ 15 The Principle of Relativity and Modern Cosmology
In special relativity, the transformations preserving this form-invariance between inertial frames are the Lorentz transformations  [^50] In the special theory of relativity the form-invariant transformations between inertial frames are Lorentz transformations [58]. For a general discussion about the Principle of Relativity we suggest reading the paper [59]. . According to the special principle of relativity, all inertial frames are equivalent concerning physical laws, but not necessarily concerning physical facts.
@@
More recent observations by astronomers also cast doubt on the CMB dipole, being the ultimate representative of the solar peculiar velocity. In recent years observations have emerged hinting at an anisotropic universe. The discovery of the preferred direction in the universe was serendipitous. Until very recently the velocity of the solar system in the rest frame of the universe is inferred from CMB temperature dipole anisotropy. Obviously, an independent measurement of this velocity is needed to fully establish the kinematical origin of the dipole. Another such quantity that could be employed to look for departures from isotropy is the angular distribution of distant radio sources in the sky. This could provide an independent check on the interpretation of CMB dipole. The radio data clearly indicate that significantly larger dipoles exist in the rest frame of the radio galaxies. While the velocity of the solar system inferred from the CMB temperature dipole anisotropy is 370 km/s, the radio dipole measurements find the speed of motion to be around 1000 km/s (i.e. to be around three times larger than that of CMB). From the Hubble diagram of quasars, the motion of the solar system is derived, which is out to be the largest value ever found, $8\cdot 10^{3}$ km/s  [^51] Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe. The various dipoles, including CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [60, 61, 62]..

### v12

@@ 12.5 Metrics in Accelerated Frames
In this section, we continue our exploration of non-reciprocity in special relativity. We begin by examining the implications of assuming that the inertial frame has no history of acceleration and that the space-time metric, as perceived by an observer at rest, is diagonal.
@@
Consider two emitters within this inertial frame: the first remains stationary, while the second moves with velocity $v$. A Galilean boost transforms the metric associated with the moving emitter, resulting in the form given in Eq. (11). In the coordinates of the inertial frame, the non-accelerated emitter is described by the standard diagonal Minkowski metric: $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$.
Now, let us shift our attention to the accelerated frame in which the second emitter is at rest, and the first appears to move with velocity v-v. This raises a fundamental question: Does the principle of reciprocity still apply in this context?
Conventional theory holds that acceleration does not intrinsically violate the symmetry of motion between inertial and non-inertial frames. Therefore, an observer in the accelerated frame should also describe the metric of the stationary (second) emitter as diagonal. By reciprocity, the moving (first) emitter should then be described in this frame by a metric analogous to Eq. (11), differing only by the sign of $v$.
However, this leads to a deeper inquiry into the principle of reciprocity itself. The principle of reciprocity states that all inertial frames are equivalent. However, this relativity principle does not hold and the Lorentz covariance of all physical equations does not dictate the reciprocal symmetry of nature.
A geometric analysis of space-time in the accelerated frame reveals this asymmetry. Specifically, within the accelerated frame, the Langevin metric—given in Eq. (15)—emerges as the correct description of the second emitter’s space-time. This metric governs electrodynamics in the accelerated frame and departs from the simple diagonal form of inertial frames.
This departure has significant consequences: Maxwell’s equations, valid in inertial frames, no longer retain their form for an observer at rest in the accelerated frame. This breakdown signals a deeper structural difference between inertial and accelerated reference frames. Notably, because we adopt an absolute time coordinatization, the initial conditions—such as the orientation of radiation wavefronts—remain identical in both frames.
We now arrive at a paradox. The equivalence of all inertial frames is often supported by the claim that, after an appropriate coordinate transformation (e.g., diagonalization of the metric tensor), both the inertial and accelerated frames yield diagonal metrics. This seems to imply symmetry: both frames appear to be related by a Lorentz transformation and should be physically equivalent.  [^46] Many physicists argue that the universality of Maxwell’s equations across inertial frames rules out the existence of any privileged frame. As Dieks notes [51], “As we already mentioned, it is a basic principle of the special theory of relativity that the line element $ds$ supplies all information about the physics of the situation, as described in the given coordinates.”
However, a key asymmetry becomes evident when we consider the physical history of the frames. While the inertial frame remains unaccelerated, the accelerated frame retains a memory of its dynamical past. This memory is encoded in the metric structure. After diagonalization, the metric tensor in the accelerated frame undergoes a discontinuous shift—from a form such as $g_{00}=(1-v^{2}/c^{2})<1,g_{01}=-v/c$ to the standard inertial values: $g_{00}=1,g_{01}=0$. This discontinuity highlights the fact that the Lorentz transformation does not preserve the smoothness of the metric across such transitions. Hence, the apparent symmetry between frames breaks down when physical considerations—specifically, the history of acceleration—are taken into account.
This leads us to a subtle but important conclusion: although the mathematical form of the metric may appear similar in both frames, the physical context—particularly the presence or absence of past acceleration—introduces observable differences. The accelerated frame thus retains a “memory” of its non-inertial history, breaking the naive equivalence suggested by purely kinematic symmetry arguments.
@@ 13.1 The Concept of Time in Special Relativity
Here is a simple example that illustrates the concept clearly. Suppose we know the law governing muon decay in its rest frame. When we apply a Lorentz transformation to this decay law, we find that, in the laboratory frame, the characteristic lifetime of the particle appears to increase from $\tau_{0}$ to $\gamma\tau_{0}$. This result can be interpreted as follows: after traveling a distance of $\gamma v\tau_{0}$, the number of muons observed in the lab frame is reduced to half of the original population. In this context, the time measured corresponds to a ”length”. This leads to an intriguing question: could this interpretation apply to all time measurements within the framework of special relativity? We argue that the answer is yes.
@@ 13.2 Interference Phenomena
Suppose the laboratory observer measures the wavelength of a standing wave. The relation $dl_{i}^{2}=(-g_{11}+g_{01}g_{01}/g_{00})dx^{2}$ describes the connection between the spatial line element and the metric. Using Eq. (11), we find $dl_{i}^{2}=dx^{2}/(1-v^{2}/c^{2})$, which defines the spatial geometry in the context of the non-diagonal metric. This result reflects the Lorentz contraction of measuring rods. When analyzing the measured data, the laboratory observer finds that the wave number is given by $\sqrt{1-v^{2}/c^{2}}\omega_{0}/[c(1-v/c)]=\omega_{0}\sqrt{1+v/c}/[c\sqrt{1-v/c}]$. Notably, this is the same factor one obtains under the assumption of a diagonal metric. In other words, the laboratory observer will measure the same two-way speed of light regardless of the choice of metric representation.
@@ 14.1 Second-Order Optical Experiments and Special Relativity
An illustrative example of time dilation is observed in the resonance absorption of gamma rays in Mössbauer rotating disk experiments. When measuring the resonance absorption for the same relative velocity between the source and the observer, a blue shift is detected when the source is positioned at the center of the disk, while a red shift appears when the source is located at the rim. This result reveals an asymmetry in the transverse Doppler effect. [^47] The first experimental confirmation of time dilation via resonance absorption of gamma rays was reported by Champeney, Isaak, and Khan in 1965 [52]. Notably, when both the source and observer are situated at the rim, no frequency shift is observed. In such a configuration, a pulse emitted by one accelerated observer on the rotating disk, which appears redshifted at the center due to time dilation, is received by a diametrically opposite observer on the rim without any frequency change. This occurs despite both observers moving at the same speed $v$ in opposite directions in the laboratory frame. Their mutual time dilation effects effectively cancel the frequency shift in this scenario.
@@ 14.3 Wigner Rotation and the Trouton-Noble Experiment
Consider two opposite point charges +e+e and e-e in the Earth-based frame; the radius-vector pointing from e-e to +e+e be denoted by $\vec{r}_{n}$. If the charges are at rest the force acting upon +e+e can be written $\vec{F}_{n}=e\vec{E}=-e^{2}\vec{r}_{n}/r^{3}_{n}$. As the force acts in the direction $\vec{r}_{n}$ the moment of force produced by the pair of charges vanishes, $\vec{M}_{n}=\vec{r}_{n}\times\vec{F}_{n}=0$.
If the pair of charges is made to move with a velocity $\vec{v}$ in the Sun-centered frame then the positive charge will be under the action of the Coulomb attraction of e-e and also under the influence of the magnetic field $\vec{B}=e(\vec{v}\times\vec{r})/(cr^{3})$. Thus the total force acting upon $e$ is given by $\vec{F}+e^{2}(\vec{v}\times(\vec{v}\times\vec{r}))/(c^{2}r^{3})$. Since $\vec{v}\times(\vec{v}\times\vec{r})=\vec{v}(\vec{v}\cdot\vec{r})-v^{2}\vec{r}$ we find that the moment of force produced by the pair of charges is equal to $\vec{M}=e^{2}(\vec{v}\cdot\vec{r})(\vec{r}\times\vec{v})/(c^{2}r^{3})$. Denoting the angle between $\vec{v}$ and $\vec{r}$ by $\theta$, we find for the absolute value of moment of force $M=e^{2}v^{2}\sin{\theta}\cos{\theta}/(c^{2}r)$.
@@
The absence of torque on the capacitor in the Earth-based frame is fully consistent with special relativity and can be explained in a straightforward manner. The electrodynamic equations in the Earth-based frame can be derived by applying a Galilean transformation (with velocity v-v) to Maxwell’s equations. Assuming the magnetic field $\vec{B}=0$ in the original frame, the transformed fields are given by $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$, where $\vec{v}$ is the velocity of the Earth relative to the Sun-based frame. Consequently, the induced magnetic field in the Earth-based frame does not exert any torque on the capacitor, which is at rest in that frame.
The presence of a magnetic field gives rise to a magnetic force, which, in turn, exerts a torque on a charge in the Sun-based inertial frame. This leads to an apparent paradox: the mechanical equations governing force and torque appear to differ between inertial frames. Specifically, a torque—and thus a time-dependent change in three-dimensional angular momentum—exists in one inertial frame but not in another that is in relative motion.  [^49] In [55], Jackson discusses the apparent paradox that different mechanical equations for force and torque govern the motion of a charged particle in different inertial frames. A similar paradox arises in the standard explanation of the Trouton–Noble experiment. In the usual approach [55], this paradox stems from the hidden assumption that all inertial frames share a common three-dimensional space. However, we demonstrate that a relativistically accurate description of the Trouton–Noble experiment can be achieved only by taking into account the Wigner rotation effect. We argue that the root of this paradox lies in a flawed assumption: that an Earth-based observer and a Sun-based observer share a common three-dimensional space. According to special relativity, observers following different worldlines (i.e., with different trajectories through spacetime) possess different three-dimensional hypersurfaces of simultaneity—different 3-spaces. The standard treatment of the Trouton-Noble experiment implicitly assumes that the coordinate axes $(x_{n},y_{n},z_{n})$ of the moving Earth-based observer are parallel to the axes $(x,y,z)$ of the stationary Sun-based observer. We demonstrate that a consistent explanation, fully aligned with special relativity, emerges only when the equations of motion are expressed in covariant form. This approach properly accounts for the relativistic transformation of coordinate systems and resolves the apparent contradiction.
@@ 14.4 Time Dilation in Moving Atomic Clock
In 1971, Hafele and Keating conducted a landmark experiment in which four atomic clocks were synchronized with a laboratory reference clock and then flown around the Earth aboard commercial aircraft—two traveling eastward and two westward.  [^50] This experiment provided the first experimental confirmation of time dilation as predicted by special relativity, using atomic clocks in circular motion. An important follow-up came from the Global Positioning System (GPS), which also involves relativistic effects. GPS satellites, numbering around 80, orbit the Earth at velocities of approximately 4 km/s. The system demonstrates that satellite clocks tick more slowly due to their relative motion, with a measured discrepancy of about 7.1 microseconds per day. Interestingly, while there is a noticeable time difference between the clocks on Earth and those in orbit, no asymmetries are observed among the satellites themselves. This suggests that relative velocity alone does not account for time dilation. Furthermore, the experiment revealed no effect of Earth’s orbital velocity on clock rates, a result attributed to the fact that both the satellites and Earth-based clocks share a common orbital path around the Sun, rendering orbital effects unobservable in this context [56, 57]. Upon return, the airborne clocks were compared with the laboratory reference clock. The results showed that the clocks carried on the planes had run slower than those left stationary on Earth. The observed discrepancies matched the predictions of special relativity, confirming the role of relative acceleration in time dilation. Notably, the experiment found no measurable influence of Earth’s orbital motion on clock rates, reinforcing the idea that time dilation effects are tied to local motion relative to the chosen inertial frame.
The null result of the experiment described above can be understood by considering that the setup involves a circular path geometry. A more careful analysis reveals that it is essential to account for the fact that the moving clocks are ultimately compared to a reference clock located at the same point and at rest relative to the Earth. The moving clock travels around the reference clock along a circular path with velocity $w$. We will analyze the effect by expanding terms up to second order in $w/c$ and $v/c$, where $v$ is the Earth’s orbital velocity. The inner rate of the moving clock depends on velocity $w$, so that $d\tau=dt_{n}[1-v^{2}/(2c^{2})-w^{2}/(2c^{2})+\vec{v}\cdot\vec{w}(t_{n})/c^{2}]$. Since the reference clock is at rest in the Earth-based frame, the reading of its proper time is $d\tau=dt_{n}(1-v^{2}/2)$. Analyzing the geometry of the situation, we see that for motion along a closed circular path, the cross term averages out over a complete cycle. Specifically, the integral of the mixed term vanishes: $\int\vec{v}\cdot\vec{w}(t_{n})dt_{n}=\int\vec{v}\cdot d\vec{s}=0$. Therefore, no observable effect from the Earth’s orbital velocity remains, which explains the null result.
@@ 15 The Principle of Relativity and Modern Cosmology
In special relativity, the transformations preserving this form-invariance between inertial frames are the Lorentz transformations.  [^51] In the special theory of relativity the form-invariant transformations between inertial frames are Lorentz transformations [58]. For a general discussion about the Principle of Relativity we suggest reading the paper [59]. According to the special principle of relativity, all inertial frames are equivalent concerning physical laws, but not necessarily concerning physical facts.
@@
More recent observations by astronomers also cast doubt on the CMB dipole, being the ultimate representative of the solar peculiar velocity. In recent years observations have emerged hinting at an anisotropic universe. The discovery of the preferred direction in the universe was serendipitous. Until very recently the velocity of the solar system in the rest frame of the universe is inferred from CMB temperature dipole anisotropy. Obviously, an independent measurement of this velocity is needed to fully establish the kinematical origin of the dipole. Another such quantity that could be employed to look for departures from isotropy is the angular distribution of distant radio sources in the sky. This could provide an independent check on the interpretation of CMB dipole. The radio data clearly indicate that significantly larger dipoles exist in the rest frame of the radio galaxies. While the velocity of the solar system inferred from the CMB temperature dipole anisotropy is 370 km/s, the radio dipole measurements find the speed of motion to be around 1000 km/s (i.e. to be around three times larger than that of CMB). From the Hubble diagram of quasars, the motion of the solar system is derived, which is out to be the largest value ever found, $8\cdot 10^{3}$ km/s.  [^52] Recent observations of the cosmic dipole anisotropies is revolutionized our understanding of the universe. The various dipoles, including CMB dipole, all pointing along the same direction, suggest a preferred direction in the universe, raising thereby uncomfortable questions about the cosmological principle, the basis of the standard model in modern cosmology [60, 61, 62].